An object, s velocity is 4m/s and the mass is 60 kg. What is the object, s kinetic energy.
1
Expert's answer
2019-07-01T09:28:57-0400

We can find the object's kinetic energy from the formula:


"KE = \dfrac{1}{2} mv^2 = \dfrac{1}{2} \cdot 60 kg \cdot (4 \dfrac{m}{s})^2 = 480 J."

Answer:

"KE = 480 J."